OpenStudy (anonymous):

If f(x) = sin(x) for all x, then the average value of f on the interval [0, π] is (A) ½ (B) 1/π (C) π/2 (D) 2/π

hartnn (hartnn):

you know we have to evaluate a integral for this, right ?

hartnn (hartnn):

\(\huge \int \limits_0^{\pi} \dfrac{\sin x }{(\pi-0)}dx \)

hartnn (hartnn):

i have used the formula for average value over [a,b] of f(x) : \(\huge \int \limits_a^b \dfrac{f(x)}{b-a}dx\)

hartnn (hartnn):

so tell me whether you got how we formed that integral :)

hartnn (hartnn):

because if you know simple integration formulas, you will be able to evaluate that integral easily.

hartnn (hartnn):

@muhammadshuaib can you tell me whether you understand how i formed that integral

OpenStudy (anonymous):

no yar i am studing fsc level..........

hartnn (hartnn):

so you have not learnt integration yet ?

OpenStudy (anonymous):

no i know about integration but not of higher level........ b/c in pak integration study not of average......in fsc level

hartnn (hartnn):

solving this integral is basic level only \(\huge \int \limits_0^{\pi} \dfrac{\sin x }{(\pi-0)}dx\) can you solve it ?

hartnn (hartnn):

i got that using the standard formula of average f(x) = sin x a= 0 b = pi

OpenStudy (anonymous):

i solve it answering pi/2 are this correct

hartnn (hartnn):

how did u get pi/2 can you show me steps ?

hartnn (hartnn):

or was that pi/2 just a guess ?

OpenStudy (anonymous):

-cos(x)/pi]limit =-{cos0-cospi}/pi =-{1-(-1)}/pi =-2/pi

hartnn (hartnn):

nice try! but we put the upper limit first!

hartnn (hartnn):

- { cos pi - cos 0}/ pi = + 2/pi which is your final answer :)

OpenStudy (anonymous):

=-{cospi-coso}/pi =-{-1-1}/pi =-[-2]/pi =2/pi

hartnn (hartnn):

correct! good :)

hartnn (hartnn):

yes ?

OpenStudy (anonymous):

your name

hartnn (hartnn):

you can call me hartnn :)

OpenStudy (anonymous):

no i asked you only name yar.......

hartnn (hartnn):

why do you need my name ?

OpenStudy (anonymous):

my name is Muhammad shuaib pakistani enshaALLah will chat you later now i m going to pray a prayer

hartnn (hartnn):

nice to meet you :) have a good day :)